OpenStudy (anonymous):

find the force of a bacteria of velocity 20nmμs-1 in the bloodstream of a patient along a vein of length 1mm, if the mass of the bacteria is o.o1μg.

OpenStudy (konradzuse):

Force is in N, do you know the formula?

OpenStudy (anonymous):

F=ma, that's all i know

OpenStudy (konradzuse):

what exactly is velocity 20nmμs-1 ???

OpenStudy (konradzuse):

@TuringTest

OpenStudy (anonymous):

its the rate at which the bacteria is travelling 20 newton metres, and the μ is the symbol for mu,,,,,i'm not really familiar with these units

OpenStudy (anonymous):

hello?????!!!!!!!!!!!!

OpenStudy (ghazi):

\[V= \frac{ 20 * 10^ {-9}}{ 10^{-6} }=20*10^{-3} m/s\] and length = 0.001 m now firstly, find the acceleration using the relation \[U^2=2*a*D\] where D is distance \[a=\frac{ U^2 }{ 2D }=(20*10^{-3})/2*10^{-3}=10 m/s^2\] now force F=mass * acceleration\[F= 10^{-8}* 10= 10^{-7} N\] = 10^(-7) N hope this helps you and remember every power is negative :D

OpenStudy (anonymous):

thank you sooooo much!!!!!!!!!!!!11

OpenStudy (ghazi):

:D YW

OpenStudy (anonymous):

what's the correct way of writing 10^(-7) N?

OpenStudy (ghazi):



OpenStudy (anonymous):

oh ok and one more thing is that a multiplication sign?

OpenStudy (ghazi):

NO, that is power of 10 or exponent of 10

OpenStudy (anonymous):

i meant this

OpenStudy (ghazi):

i dont know what you meant

OpenStudy (anonymous):

sorry i tried to draw i meant the sign between the 20 and 10(-7) at the first part where you calculated velocity

OpenStudy (ghazi):

thats multiplication hope you have seen it for multiplication thing

OpenStudy (anonymous):

oh ok thanks for you help! my teacher is a hand full he's so unfair!

OpenStudy (ghazi):

its okay :)

OpenStudy (anonymous):

ok take care bye! i now know who to come for when i need help

OpenStudy (ghazi):

see ya :)